OpenStudy (anonymous):

Find the value of f(-4) if f(x) = x^2 + 7.

Parth (parthkohli):

\[f(x) = x^2 + 7 \implies f(y) = y^2 + 7\implies f(z) = z^2 + 7\implies \cdots\]\[\implies f(-4) = (-4)^2 + 7\]

OpenStudy (anonymous):

@cyprus23 you keep asking the same "plug and play" stuff. Why don't show us how do you do it.